OpenStudy (anonymous):

What is the derivative of 5xsecx+x^3cosx+17(pi)?

myininaya (myininaya):

You will need to know the derivative of a constant, the derivative of cos(x),the derivative of sec(x), and the derivative of x, the product rule, and the power rule. So let's look at each term individually: 5xsec(x) x^3 cos(x) 17pi

myininaya (myininaya):

5xsec(x) is a product [(5x) * (sec(x))]'=? You will need to know the product rule. Hint here: (sec(x))'=tan(x)sec(x) --- [x^3]'=? You will need to know the power rule. ---- (cos(x))'=-sin(x) ------ 17pi is a constant so (17pi)'=? What does the constant rule say?

OpenStudy (anonymous):

[x^3]'= \[3x ^{2}\]

myininaya (myininaya):

yes

OpenStudy (anonymous):

will the derivative of 17pi be 0?

myininaya (myininaya):

yep

myininaya (myininaya):

So the last part to look at is (5xsec(x))'

myininaya (myininaya):

(5x)'=?

myininaya (myininaya):

Also the product rule is (u*v)'=u'*v+u*v'

OpenStudy (anonymous):

(5x)'= x ?

myininaya (myininaya):

That is like saying the derivative of x is x which is not.

myininaya (myininaya):

The slope of y=5x is 5 so the derivative of (5x) is 5.

OpenStudy (anonymous):

ooh okay so (5xsec(x))' = 5tan(x)sec(x) ?

myininaya (myininaya):

Well that is a product you need to use the product rule not a madeup one.

myininaya (myininaya):

(5xsec(x))' = (5x)'sec(x)+5x(sec(x))' by product rule

OpenStudy (anonymous):

what about tan(x) ? All of my answer choices have tan in them?

myininaya (myininaya):

Right which is a good thing because our answer as tan in it.

myininaya (myininaya):

has*

OpenStudy (anonymous):

based on what we have said I am leading towards answer choice C which is 5sec(x)tan(x)+3x^2sin(x) is that the answer?

myininaya (myininaya):

So...well again (5xsec(x))' is equal to (5x)'sec(x)+5x(sec(x))'

myininaya (myininaya):

You know (5x)' and (sec(x))' Just replace them in this: (5x)'sec(x)+5x(sec(x))'

OpenStudy (anonymous):

so 5secx+5xsecxtanx?

OpenStudy (anonymous):

cause my other answer choice is 5secx+5xsecxtanx+3x^2cosx-x^3sinx , but I am not sure how they got x^3sinx

myininaya (myininaya):

What is the derivative of 5xsecx+x^3 + cosx+17(pi) this was the question right? this is what you wrote

myininaya (myininaya):

You didn't write x^3cos(x) you wrote x^3+cos(x)

OpenStudy (anonymous):

:O I meant x^3cos(x)

myininaya (myininaya):

Use the product rule again since we have a product.

OpenStudy (anonymous):

i think the answer is 5sec(x)tan(x)+3x^2sin(x)

myininaya (myininaya):

What about the other parts?

myininaya (myininaya):

We just said (5xsec(x))'=5sec(x)+5xsec(x)tan(x) Now how do we find the derivative of x^3cos(x)?

OpenStudy (anonymous):

well 17pi =0 right?

myininaya (myininaya):

(17pi)'=0

myininaya (myininaya):

We need to find (x^3cos(x))' Again we need to use the product rule (x^3cos(x))' = (x^3)'cos(x)+x^3(cos(x))'

OpenStudy (anonymous):

=-3x^2sinx

myininaya (myininaya):

nope there are two terms

OpenStudy (anonymous):

yes well that 2 times: -3x^2sinx+3x^2sinx

myininaya (myininaya):

no no...

myininaya (myininaya):

(x^3)'=3x^2 (cos(x))'=-sin(x) (x^3cos(x))' WE are finding the derivative of a product so we must use the product rule (x^3)'cos(x)+x^3(cos(x))' Now looking above 3x^2cos(x)+x^3(-sin(x))

OpenStudy (anonymous):

so 3x^2cosx-x^3sinx?

myininaya (myininaya):

for that part yes

OpenStudy (anonymous):

so the answer is 5secx+5xsecxtanx+3x^2cosx-x^3sinx

myininaya (myininaya):

yes

OpenStudy (anonymous):

thank you!

myininaya (myininaya):

np I hope you understand the product rule now